Use bond energies from this table to estimate the heat of reaction for the following process:4. Use bond energies from this table to estimate the heat of reaction for the following process: HH н C= C + Cl - H- C- C-H CSome Bond Energies (kJ/mol) Н-Н O=O • Diatomic molecules (actual values) 432.0 Cl-ci 240.0 C=0 1071 493.6 Other bonds (avg. v
Answer #1

On reactant side we have:

4 C-H bond

1 C=C bond

1 Cl-Cl bond

On product side we have:

4 C-H bond

1 C-C bond

2 C-Cl bond

Use:

Delta H = BE(reactant) - BE(product)

= 4*BE(C-H) + 1*BE(C=C) + 1*BE(Cl-Cl) - 4*BE(C-H) - 2*BE(C-Cl) - 1*BE(C-C)

= 1*BE(C=C) + 1*BE(Cl-Cl)- 2*BE(C-Cl) - 1*BE(C-C)

= 1*615 + 1*240 - 2*330 - 1*345

= -150 KJ/mol

Answer: -150 KJ/mol